Consider the following term structure of interest rates:
• 6 months – 1.0%
• 1 year – 1.1 %
• 1.5 years – 1.4%
• 2 years – 1.8%
• 2.5 years – 2.4%
Find the price of a 4% bond paying coupons semi-annually and that matures in 1.5 years.
The bond has par (face) value of $1,000.
Answer #1

Face value or Par value of the bond = $1000

C = coupon payment = 4% of $1000 = $40 but semiannual coupon, therefore C = $40/2 = $20

n = number of payments (time remaining to maturity) = 1.5 years; therefore number of payments n = 1.5 *2 = 3

The cash flows of bond is following -

Payments

1

2

3

Cash Flows

$20

$20

$1,020

The 6-months rate is 1.0% APR, compounded semi-annually, therefore 1.0%/2 = 0.5% semiannual rate

The 1.0-year rate is 1.1% APR, compounded semi-annually, therefore 1.1%/2 = 0.55% semiannual rate

The 1.5-year rate is 1.4% APR, compounded semi-annually, therefore 1.4%/2 = 0.7% semiannual rate

Price of the bond = present value of cash flows of bond

= $20/ (1+0.5%) ^1 + $20/ (1+0.55%) ^2 + $1,020/ (1+ 0.7%) ^3

= $19.90 + $19.78 + $998.88

= $1038.56

Price of the bond is $1,038.56